OpenStudy (anonymous):

Simplify the given expression. 1/6(3x+5y)+2/3(3/5x-6y)

OpenStudy (midhun.madhu1987):

can u draw or put the screenshot of the question?

OpenStudy (dangerousjesse):

You don't have to draw it, just use the equation option. Would probably be easier :P

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }\left( 3x+5y \right)+\frac{ 2 }{ 3 }\left( \frac{ 3 }{ 5 }x-6y \right)\]

OpenStudy (isaiah.feynman):

Distribute the 1/6 and 2/3 first.

OpenStudy (dangerousjesse):

Cancel common terms in the numerator and denominator of \(\frac{3 (x-6 y)×2}{3×5}\). \(\frac{3 (x-6 y)×2}{3×5} = \frac{3}{3}×\frac{(x-6 y)×2}{5} = \frac{(x-6 y)×2}{5}\): \(\frac{3 x+5 y}{6}+\frac{2 (x-6 y)}{5}\) Put the fractions in \(\frac{3 x+5 y}{6}+\frac{(x-6 y)×2}{5}\) over a common denominator. Put each term in \(\frac{3 x+5 y}{6}+\frac{(x-6 y)×2}{5}\) over the common denominator 30: \(\frac{3 x+5 y}{6}+\frac{(x-6 y)×2}{5} = \frac{5 (3 x+5 y)}{30}+\frac{12 (x-6 y)}{30}:\) \(\frac{5 (3 x+5 y)}{30}+\frac{12 (x-6 y)}{30}\) Combine \(\frac{5 (3 x+5 y)}{30}+\frac{12 (x-6 y)}{30}\) into a single fraction. \(\frac{5 (3 x+5 y)}{30}+\frac{12 (x-6 y)}{30} = \frac{5 (3 x+5 y)+12 (x-6 y)}{30:}\) \(\frac{5 (3 x+5 y)+12 (x-6 y)}{30}\) Distribute 5 over \(3 x+5 y.\) \(5 (3 x+5 y) = 15 x+25 y:\) \(\frac{15 x+25 y+12 (x-6 y)}{30}\) Distribute 12 over \(x-6 y.\) \(12 (x-6 y) = 12 x-72 y:\) \(\frac{15 x+25 y+12 x-72 y}{30}\) Group like terms in \(15 x+25 y+12 x-72 y.\) Grouping like terms, \(15 x+25 y+12 x-72 y = (15 x+12 x)+(25 y-72 y):\) \(\frac{(15 x+12 x)+(25 y-72 y)}{30}\) Combine like terms in \(25 y-72 y.\) \(25 y-72 y = -47 y:\) \(\frac{-47 y+(15 x+12 x)}{30}\) Add like terms in 15 x+12 x. \(15 x+12 x = 27 x:\)

OpenStudy (dangerousjesse):

Refresh the page if you can't see certain signs.

OpenStudy (anonymous):

a. 3x + 4y b. \[\frac{ 9 }{ 10 }x-\frac{ 19 }{ 6}y\] c. \[\frac{ 2 }{ 3 }y+14x\] d. 32x+46y